### v3

# Deep Koopman Learning of Nonlinear Time-Varying Systems

Abstract
This paper presents a data-driven approach to approximate the dynamics of a nonlinear time-varying system (NTVS) by a linear time-varying system (LTVS), which is resulted from the Koopman operator and deep neural networks. Analysis of the approximation error between states of the NTVS and the resulting LTVS is presented. Simulations on a representative NTVS show that the proposed method achieves small approximation errors, even when the system changes rapidly. Furthermore, simulations in an example of quadcopters demonstrate the computational efficiency of the proposed approach.
1 Introduction
In recent years, data-driven methods have received a significant amount of research attention due to the increasing complexity of the autonomous systems in both dynamics [1, 2, 3, 4] and mission objectives [5, 6, 7]. In the direction of learning system dynamics, the Koopman operator has recently been proven to be an effective method to approximate a nonlinear system by a linear time-varying system based on state-control pairs [2, 3, 4]. Along this direction, two popular methods dynamic mode decomposition (DMD) and extending dynamic mode decomposition (EDMD) are used to lift the state space to a higher-dimensional space, where the evolution is approximately linear [8]. However, choosing the proper observable functions for the lifting transformation is still an open question, and the potentially large lifted dimension may hinder real-time applications.
Recent work has proposed several methods for choosing proper observable functions of Koopman-based methods for time-invariant systems. Lusch et al. [9] proposed applying deep learning methods to discover the eigenfunctions of the approximated Koopman operator. Yeung et al. [10, 11, 12, 13] introduced deep neural networks (DNN) as observable functions of the Koopman operator, which are tuned based on collected state-control pairs by minimizing a properly defined loss function. While some work, such as [14], has extended the DMD method to approximate nonlinear time-varying systems (NTVS) that change sufficiently slowly by linear time-varying systems (LTVS), this method is not directly applicable to approximate nonlinear systems with rapidly changing dynamics.
In this paper, we propose a deep Koopman learning method to approximate NTVS, which employs DNN as the observable function of the Koopman operator and adjusts both the DNN and the approximated dynamical system simultaneously. This is achieved by tuning the DNN parameters based on the latest state-control data pairs to track the unknown NTVS.
Compared to existing results in [14], the proposed method is able to approximate an NTVS which does not necessarily change slowly. Contributions of this paper are summarized as follows:
  • We propose a deep Koopman representation formulation for the NTVS and provide a practical online algorithm for implementation.
  • We investigate the error bound of the system state estimation of the proposed method.
  • We perform a convergence analysis of the proposed method concerning the observable function of DNN.
This paper is organized as follows. In Section 2, we state the problem. Section 3 presents the main results. The numerical simulations are exhibited in Section 4. Finally, Section 5 concludes the paper.
[…]
2 Problem Formulation
Consider an NTVS, the dynamics of which is unknown. Let $x_{t}\in\mathbb{R}^{n}$ and $u_{t}\in\mathbb{R}^{m}$ denote its state and control input at time $t$, respectively. $t\in[0,\infty)$ denotes the continuous-time index.
Suppose the states and control inputs can be obtained from unknown continuous NTVS at certain sampling time instances $t_{k}\in[0,\infty)$ with $k=0,1,2,\cdots$ the index of sampled data points. For notation brevity, one denotes $x_{k}\coloneqq x_{t_{k}}$ and $u_{k}\coloneqq u_{t_{k}}$ as the $k$-th observed system state and control input, respectively, in the remainder of this manuscript. Then one can partition the observed states-inputs pairs as the following series of data batches
$\mathcal{B}_{\tau}=\{x_{k},u_{k}:k\in\mathbb{K}_{\tau}\},\quad\tau=0,1,2,\cdots,$ (1)
where
$\mathbb{K}_{\tau}=\{k_{\tau},k_{\tau}+1,k_{\tau}+2,\cdots,k_{\tau}+\beta_{\tau}\}$
denotes the ordered labels set of sampling instances for the $\tau$-th data batch $\mathcal{B}_{\tau}$ with $\beta_{\tau}$ positive integers such that
$k_{\tau}=\sum_{i=0}^{\tau-1}\beta_{i},\quad\tau\geq 1,\quad k_{0}=0.$
It follows that
$k_{\tau+1}=k_{\tau}+\beta_{\tau},\quad\tau=0,1,2,\cdots,$
which implies the last data in the $\tau$-th data batch is the first data in $(\tau+1)$-th data batch. For notation simplicity, one defines $\mathcal{B}_{\tau}^{x}\coloneqq\{x_{k}:k\in\mathbb{K}_{\tau}\}$, $\mathcal{B}_{\tau}^{u}\coloneqq\{u_{k}:k\in\mathbb{K}_{\tau}\}$ in the rest of this manuscript. An illustration of the above indexes is shown in Fig. 1.
Figure 1: Time indexes, where blue color denotes $k_{\tau}$.
This paper aims to develop an iterative method that approximates the dynamics of an unknown NTVS based on available data batches $\mathcal{B}_{\tau}$ by a linear time-varying discrete-time system. One way to achieve such a linear approximation is by employing the Koopman operator as in [10, 11, 12, 13]. Namely, based on the data batch $\mathcal{B}_{\tau}$, one finds a nonlinear mapping $g(\cdot,\theta_{\tau}):\mathbb{R}^{n}\rightarrow\mathbb{R}^{r}$ parameterized by $\theta_{\tau}\in\mathbb{R}^{q}$ 111 Here, $g(\cdot,\theta_{\tau})$ is usually represented by a DNN with a known structure $g$ and an adjustable parameter $\theta_{\tau}\in\mathbb{R}^{q}$. and constant matrices $A_{\tau}\in\mathbb{R}^{r\times r}$, $B_{\tau}\in\mathbb{R}^{r\times m}$, $C_{\tau}\in\mathbb{R}^{n\times r}$ such that for $k\in\mathbb{K}_{\tau},k<k_{\tau}+\beta_{\tau}$, the following holds approximately,
$g(x_{k+1},\theta_{\tau})$ $=A_{\tau}g(x_{k},\theta_{\tau})+B_{\tau}u_{k},$ (2)
$x_{k+1}$ $=C_{\tau}g(x_{k+1},\theta_{\tau}).$ (3)
Here, $g(\cdot,\theta_{\tau})$, $A_{\tau}$, $B_{\tau}$, $C_{\tau}$ achieved from $\mathcal{B}_{\tau}$ are put together in the following $\mathcal{K}_{\mathcal{B}_{\tau}}$:
$\mathcal{K}_{\mathcal{B}_{\tau}}\coloneqq\{g(\cdot,\theta_{\tau}),A_{\tau},B_{\tau},C_{\tau}\},$ (4)
which is called a deep Koopman representation (DKR) in this manuscript.
Based on the DKR in (4), one could introduce $\hat{x}_{k}\in\mathbb{R}^{n}$ for $k\in\mathbb{K}_{\tau},k<k_{\tau}+\beta_{\tau}$ as follows:
$g(\hat{x}_{k+1},\theta_{\tau})$ $=A_{\tau}g(\hat{x}_{k},\theta_{\tau})+B_{\tau}\hat{u}_{k},$ (5)
$\hat{x}_{k+1}$ $=C_{\tau}g(\hat{x}_{k+1},\theta_{\tau}),$ (6)
where
$\hat{u}_{k}=u_{k},\forall k\in\mathbb{K}_{\tau},k<k_{\tau}+\beta_{\tau},\quad\hat{x}_{k_{\tau}}=x_{k_{\tau}}.$ (7)
This leads to the following linear system
$\hat{x}_{k+1}=\hat{A}_{\tau}\hat{x}_{k}+\hat{B}_{\tau}\hat{u}_{k},\quad k\in\mathbb{K}_{\tau},k<k_{\tau}+\beta_{\tau},$ (8)
with $\hat{A}_{\tau}=C_{\tau}A_{\tau}C_{\tau}^{\dagger}$, $\hat{B}_{\tau}=C_{\tau}B_{\tau}$ and initial conditions in (7). The system (8) can be viewed as a linear approximation to the NTVS based on the data batch $\mathcal{B}_{\tau}$. Note that when for any data batch $\mathcal{B}_{\tau}$, one has $A_{\tau}$, $B_{\tau}$,$C_{\tau}$ remain constant for $k\in\mathbb{K}_{\tau}$. It follows that (8) is a linear time-invariant system for $k\in\mathbb{K}_{\tau}$.
To sum up, the problem of interest is to develop an iterative update rule to achieve a DKR in (4) based on data batch $\mathcal{B}_{\tau}$ in (1) such that the linear system (8) is a nice approximation of the unknown NTVS, i.e. $\hat{x}_{k}$ (8) is close to $x_{k}$ observed in $\mathcal{B}_{\tau}$ from the unknown NTVS in the sense that for any given accuracy $\varepsilon\geq 0$, $\hat{u}_{k}=u_{k}$ and $\hat{x}_{k_{\tau}}=x_{k_{\tau}}$, the estimation error $\parallel\hat{x}_{k}-x_{k}\parallel\leq\varepsilon$.
3 Main Results
This section proposes an algorithm to achieve a deep Koopman representation (DKR) that can approximate an unknown NTVS. We then investigate the estimation error between the state obtained from this DKR, as given in (8), and the observed state of the unknown NTVS.
3.1 Key Idea
Motivated by deep Koopman operator-based methods developed in [10, 11, 12, 13], an optimal $\theta_{\tau}$ for the deep Koopman representation (DKR), denoted by $\theta_{\tau}^{*}$, can be obtained by solving the following optimization problem based on the data batch $\mathcal{B}_{\tau}$:
$\theta_{\tau}^{*}=\arg\min_{\theta_{\tau}\in\mathbb{R}^{q}}\{w{\mathbf{L}}_{1}(A_{\tau},B_{\tau},\theta_{\tau})+(1-w){\mathbf{L}}_{2}(C_{\tau},\theta_{\tau})\},$ (9)
where
${\mathbf{L}}_{1}(A_{\tau},B_{\tau},\theta_{\tau})=\frac{1}{\beta_{\tau}}\sum_{k=k_{\tau}}^{k_{\tau}+\beta_{\tau}-1}\parallel g(x_{k+1},\theta_{\tau})-(A_{\tau}g(x_{k},\theta_{\tau})+B_{\tau}u_{k})\parallel^{2}$ (10)
and
${\mathbf{L}}_{2}(C_{\tau},\theta_{\tau})=\frac{1}{\beta_{\tau}}\sum_{k=k_{\tau}}^{k_{\tau}+\beta_{\tau}-1}\parallel x_{k}-C_{\tau}g(x_{k},\theta_{\tau})\parallel^{2}.$ (11)
The objectives of (10) and (11) are to approximate (2) and (3), respectively. Here, $10<w<$ 1 is a constant that combines the objective of minimizing ${\mathbf{L}}_{1}$ and ${\mathbf{L}}_{2}$. In simple terms, ${\mathbf{L}}_{1}$ and ${\mathbf{L}}_{2}$ measure the simulation errors in the lifted and original coordinates, respectively.
To solve (9), one needs to rewrite the available data batch and objective functions ${\mathbf{L}}_{1}$ and ${\mathbf{L}}_{2}$ in compact forms. Toward this end, the following notation is introduced:
${\mathbf{X}}_{\tau}$ $=[x_{k_{\tau}},x_{k_{\tau}+1},\cdots,x_{k_{\tau}+\beta_{\tau}-1}]\in\mathbb{R}^{n\times\beta_{\tau}},$
$\bar{{\mathbf{X}}}_{\tau}$ $=[x_{k_{\tau}+1},x_{k_{\tau}+2},\cdots,x_{k_{\tau}+\beta_{\tau}}]\in\mathbb{R}^{n\times\beta_{\tau}},$
$\mathbf{U}_{\tau}$ $=[u_{k_{\tau}},u_{k_{\tau}+1},\cdots,u_{k_{\tau}+\beta_{\tau}-1}]\in\mathbb{R}^{m\times\beta_{\tau}}.$
Then ${\mathbf{L}}_{1}$ in (10) and ${\mathbf{L}}_{2}$ in (11) can be rewritten as
${\mathbf{L}}_{1}=\frac{1}{\beta_{\tau}}\parallel\bar{{\mathbf{G}}}_{\tau}-(A_{\tau}{\mathbf{G}}_{\tau}+B_{\tau}\mathbf{U}_{\tau})\parallel_{F}^{2}$ (12)
and
${\mathbf{L}}_{2}=\frac{1}{\beta_{\tau}}\parallel{\mathbf{X}}_{\tau}-C_{\tau}{\mathbf{G}}_{\tau}\parallel_{F}^{2},$ (13)
where
${\mathbf{G}}_{\tau}$ $=[g(x_{k_{\tau}},\theta_{\tau}),\cdots,g(x_{k_{\tau}+\beta_{\tau}-1},\theta_{\tau})]\in\mathbb{R}^{r\times\beta_{\tau}},$ (14)
$\bar{{\mathbf{G}}}_{\tau}$ $=[g(x_{k_{\tau}+1},\theta_{\tau}),\cdots,g(x_{k_{\tau}+\beta_{\tau}},\theta_{\tau})]\in\mathbb{R}^{r\times\beta_{\tau}}.$
By minimizing ${\mathbf{L}}_{1}$ with respect to $A_{\tau},B_{\tau}$ in (12) and minimizing ${\mathbf{L}}_{2}$ regarding $C_{\tau}$ in (13), $A_{\tau},B_{\tau},C_{\tau}$ can be determined by $\theta_{\tau}$ as follows:
$[A_{\tau}^{\theta},B_{\tau}^{\theta}]$ $=\bar{{\mathbf{G}}}_{\tau}\begin{bmatrix}{\mathbf{G}}_{\tau}\\ \mathbf{U}_{\tau}\end{bmatrix}^{\dagger},$ (17)
$C_{\tau}^{\theta}$ $={\mathbf{X}}_{\tau}{\mathbf{G}}_{\tau}^{\dagger}.$ (18)
Replacing $A_{\tau}$ and $B_{\tau}$ in (10) by (17) and $C_{\tau}$ in (11) by (18), the objective function in (9) can be reformulated as
${\mathbf{L}}(\theta_{\tau})=\frac{1}{\beta_{\tau}}\!\!\sum_{k=k_{\tau}}^{k_{\tau}+\beta_{\tau}-1}\!\!\parallel\!\!\begin{bmatrix}g(x_{k+1},\theta_{\tau})\\ {x_{k}}\end{bmatrix}\!-\!K_{\tau}^{\theta}\begin{bmatrix}g(x_{k},\theta_{\tau})\\ {u_{k}}\end{bmatrix}\!\!\parallel^{2},$ (19)
with
$\ K_{\tau}^{\theta}=\begin{bmatrix}A_{\tau}^{\theta}&B_{\tau}^{\theta}\\ C_{\tau}^{\theta}&\mathbf{0}_{n\times m}\end{bmatrix}.$
Applying the existing deep Koopman operator methods developed in [10, 11, 12, 13] to achieve the DKR by solving (9) based on each $\mathcal{B}_{\tau}$ available has two shortcomings. First, computing the pseudo-inverse in (17) and (18) repeatedly while solving (9) becomes computationally expensive as $\tau$ increases. Second, $\theta_{\tau}$ must be initialized for each $\mathcal{B}_{\tau}$, which can be challenging in time-varying systems applications. To overcome these two limitations, one can apply the deep Koopman operator method to approximate the unknown NTVS efficiently, and we propose the following method.
3.2 Algorithm
Before proceeding, we need the following assumption.
Assumption 1
The matrix ${\mathbf{G}}_{\tau}\in\mathbb{R}^{r\times\beta_{\tau}}$ in (14) and $\begin{bmatrix}{\mathbf{G}}_{\tau}\\ \mathbf{U}_{\tau}\end{bmatrix}\in\mathbb{R}^{(r+m)\times\beta_{\tau}}$ are of full row rank.
Remark 1
Assumption 1 is to ensure the matrices ${\mathbf{G}}_{\tau}\in\mathbb{R}^{r\times\beta_{\tau}}$ and $\begin{bmatrix}{\mathbf{G}}_{\tau}\\ \mathbf{U}_{\tau}\end{bmatrix}\in\mathbb{R}^{(r+m)\times\beta_{\tau}}$ invertible and it naturally requires $\beta_{\tau}\geq r+m$.
Lemma 1
Given $\mathcal{K}_{\mathcal{B}_{\tau}}$ in (4), if Assumption 1 holds, then the matrices $A_{\tau+1}^{\theta},B_{\tau+1}^{\theta}$, $C_{\tau+1}^{\theta}$ can be achieved by
$=(\bar{{\mathbf{G}}}_{\tau+1}-[A_{\tau}^{\theta},B_{\tau}^{\theta}]\chi_{\tau+1})\lambda_{\tau}\chi_{\tau+1}^{T}(\chi_{\tau}\chi_{\tau}^{T})^{-1}+[A_{\tau}^{\theta},B_{\tau}^{\theta}],$ (20)
$C_{\tau+1}^{\theta}=({\mathbf{X}}_{\tau+1}-C_{\tau}^{\theta}{\mathbf{G}}_{\tau+1})\bar{\lambda}_{\tau}{\mathbf{G}}_{\tau+1}^{T}({\mathbf{G}}_{\tau}{\mathbf{G}}_{\tau}^{T})^{-1}+C_{\tau}^{\theta},$ (21)
where $\chi_{\tau}=\begin{bmatrix}{\mathbf{G}}_{\tau}\\ \mathbf{U}_{\tau}\end{bmatrix}\in\mathbb{R}^{(r+m)\times\beta_{\tau}}$, $\lambda_{\tau}=({\mathbf{I}}_{\beta_{\tau+1}}+\chi_{\tau+1}^{T}(\chi_{\tau}\chi_{\tau}^{T})^{-1}\chi_{\tau+1})^{-1}\in\mathbb{R}^{\beta_{\tau+1}\times\beta_{\tau+1}}$, $\bar{\lambda}_{\tau}=({\mathbf{I}}_{\beta_{\tau+1}}+{\mathbf{G}}_{\tau+1}^{T}({\mathbf{G}}_{\tau}{\mathbf{G}}_{\tau}^{T})^{-1}{\mathbf{G}}_{\tau+1})^{-1}\in\mathbb{R}^{\beta_{\tau+1}\times\beta_{\tau+1}}$ with ${\mathbf{G}}_{\tau+1}\in\mathbb{R}^{r\times\beta_{\tau+1}},\bar{{\mathbf{G}}}_{\tau+1}\in\mathbb{R}^{r\times\beta_{\tau+1}}$ defined in (14).
The proof of Lemma 1 will be given in [15].
To initialize the algorithm, one first needs to build the DNN $g(\cdot,\theta_{\tau}):\mathbb{R}^{n}\rightarrow\mathbb{R}^{r}$ with non-zero $\theta_{\tau}\in\mathbb{R}^{q}$. Then the matrices $A_{\tau}^{\theta}\in\mathbb{R}^{r\times r},B_{\tau}^{\theta}\in\mathbb{R}^{r\times m}$ and $C_{\tau}^{\theta}\in\mathbb{R}^{n\times r}$ can be found by solving (17) and (18), respectively, based on $\mathcal{B}_{\tau}$. When the data batch $\mathcal{B}_{\tau+1}$ becomes available, one can directly update the ${\mathbf{G}}_{\tau+1}$ and $\bar{{\mathbf{G}}}_{\tau+1}$ by following (14), which leads to the $\chi_{\tau+1}$, $\lambda_{\tau}$ and $\bar{\lambda}_{\tau}$ in Lemma 1. Thus one can update $A_{\tau}^{\theta}$, $B_{\tau}^{\theta}$ and $C_{\tau}^{\theta}$ efficiently by computing (20)-(21) instead of solving (17)-(18) repeatedly as a consequence of applying Lemma 1. Finally, an optimal $\theta_{\tau+1}^{*}$ is obtained by solving (9) with $A_{\tau+1}^{\theta},B_{\tau+1}^{\theta}$ subject to (20) and $C_{\tau+1}^{\theta}$ satisfying (21) based on $\mathcal{B}_{\tau+1}$
To sum up, we have the following algorithm, which is referred to as deep Koopman learning for time-varying systems (DKTV) in the remainder of this manuscript and its pseudocode is provided in [15].
  1. 1.
    Initialization: Build $g(\cdot,\theta_{\tau}):\mathbb{R}^{n}\rightarrow\mathbb{R}^{r}$ with $\theta_{\tau}\in\mathbb{R}^{q},\theta_{\tau}\neq\mathbf{0}_{q}$; initialize $A_{\tau}^{\theta},B_{\tau}^{\theta}$ and $C_{\tau}^{\theta}$ by solving (17) and (18), respectively, based on $\mathcal{B}_{\tau}$.
  2. 2.
    When $\mathcal{B}_{\tau+1}$ becomes available, update $A_{\tau}^{\theta}$, $B_{\tau}^{\theta}$ and $C_{\tau}^{\theta}$ according to (20) and (21), respectively.
  3. 3.
    Solve (9) to find $\theta_{\tau+1}^{*},A_{\tau+1}^{\theta^{*}},B_{\tau+1}^{\theta^{*}}$ and $C_{\tau+1}^{\theta^{*}}$.
  4. 4.
    Repeat steps $2$-$3$ as the unknown NTVS evolves.
3.3 Analysis of the Approximation Error
With a slight abuse of notation, suppose one observes the latest $\mathcal{B}_{\tau}$ with $\{x_{k-1},u_{k-1}\}$ its latest data point (i.e., $x_{k-1}\coloneqq x_{k_{\tau}+\beta_{\tau}}$, $u_{k-1}\coloneqq u_{k_{\tau}+\beta_{\tau}}$), in this subsection, we investigate the estimation errors induced by the proposed algorithm described as:
$e_{k}=\hat{x}_{k}-x_{k},$ (22)
where $\hat{x}_{k}\in\mathbb{R}^{n}$ denotes the estimated state achieved by the proposed method and $x_{k}\in\mathbb{R}^{n}$ denotes the state of the unknown NTVS evolving from $x_{k-1}\in\mathbb{R}^{n}$ and $u_{k-1}\in\mathbb{R}^{m}$. Here, to analyze the (22) with respect to $\theta_{\tau}\in\mathbb{R}^{q}$, one rewrites (5)-(6) as:
$\hat{x}_{k}=C_{\tau}(A_{\tau}g(\hat{x}_{k-1},\theta_{\tau})+B_{\tau}\hat{u}_{k-1}),$ (23)
with condition (7) hold.
Before we present the results, the following concepts about operators are introduced.
3.3.1 Preliminaries
Definition 1 (Koopman operator with control)
Consider the Hilbert space $\mathcal{F}$ and discrete-time nonlinear time-varying system $x_{k+1}=f(x_{k},u_{k},k)$ starting from time $k_{0}$, where $k\in\mathbb{Z}$ denotes its time index; $x_{k}\in\mathcal{M}$ and $u_{k}\in\mathcal{U}$ denote the system state and control input at time $k$, respectively; $\mathcal{M}\subseteq\mathbb{R}^{n}$ and $\mathcal{U}\subseteq\mathbb{R}^{m}$ denote the state space and control input space, respectively. Then the dynamics of the states of the augmented control system $z_{k},z_{k+1}$ is described by
$z_{k+1}=F(z_{k},k)\coloneqq\begin{bmatrix}f(x_{k},{\boldsymbol{u}}(0),k)\\ \mathcal{S}{\boldsymbol{u}}\end{bmatrix},$
where $z_{k}=\begin{bmatrix}x_{k}\\ \boldsymbol{u}\end{bmatrix}$ denotes the augmented control system states; $\mathcal{S}\boldsymbol{u}$ denotes the left shifting of the control sequence $\boldsymbol{u}$, i.e., $(\mathcal{S}{\boldsymbol{u}})(i)={\boldsymbol{u}}(i+1)$ and ${\boldsymbol{u}}(i)$ denotes the $i$th element of the control sequence $\boldsymbol{u}$.
We then extend the definition of the Koopman operator from [16] to discrete-time NTVS. Let operator $\mathcal{K}^{(k_{0},k)}:\mathcal{F}\rightarrow\mathcal{F}$ act on functions of state space ϕ()italic-$\phi(\cdot)\in\mathcal{F}$ with $\phi(\cdot):\mathcal{M}\times\mathcal{U}\rightarrow\mathbb{C}$ and defined with two parameters $(k_{0},k)$ as
$\mathcal{K}^{(k_{0},k)}[\phi(z_{k})]=\phi(F(z_{k},k)).$
Definition 2 ($L_{2}$-projection [17])
For brevity, we recall the definitions of $\mathcal{F}$, $\mathcal{M}$, and ϕ()italic-$\phi(\cdot)$ from Definition 1. Let $\nu\in\mathbb{R}$ be the positive measurement on $\mathcal{M}$ and assume that $\mathcal{F}=L_{2}(\nu)$. Given a set of linearly independent $\phi_{i}\in\mathcal{F}$ with $i=1,2,\cdots,r$, and define
[…]
$\Phi_{r}$ $=[\phi_{1},\phi_{2},\cdots,\phi_{r}]^{T}.$
Then for all nonzero $c\in\mathbb{R}^{r}$, the $L_{2}(\nu)$ projection of a function $\psi\in L_{2}(\nu)$ on $\mathcal{F}_{r}$ is defined as
[…]
Definition 3 (Bounded operator)
Given a Hilbert space $\mathcal{F}$, an operator $\mathcal{O}:\mathcal{F}\rightarrow\mathcal{F}$ is bounded on $\mathcal{F}$ if
$\parallel\mathcal{O}\parallel\coloneqq\sup_{f\in\mathcal{F},\parallel f\parallel=1}\parallel\mathcal{O}f\parallel<\infty.$
$\parallel\mathcal{O}\parallel$ is referred as the operator norm of $\mathcal{O}$ in the remainder of this manuscript.
Definition 4 (Strong Convergence of operator)
A sequence of bounded operators $\mathcal{O}_{i}:\mathcal{F}\rightarrow\mathcal{F}$ defined on a Hilbert space $\mathcal{F}$ converges strongly (or in the strong operator topology) to an operator $\mathcal{O}:\mathcal{F}\rightarrow\mathcal{F}$ if
$\lim_{i\rightarrow\infty}\parallel\mathcal{O}_{i}f-\mathcal{O}f\parallel=0$
for all $f\in\mathcal{F}$.
3.3.2 Analysis
We make the following assumptions to show that the error $e_{k}$ in (22) is bounded.
Assumption 2
For any $\mathcal{B}_{\tau}$ in (1), let $\{x_{i},u_{i}\}$ denote its $i$th data point observed at time $t_{i}$. The observation interval $\Delta t=t_{i+1}-t_{i}$ is sufficiently small such that for some $\mu_{x}\geq 0,\mu_{u}\geq 0$, $\parallel x_{i+1}-x_{i}\parallel\leq\mu_{x}<\infty$ and $\parallel u_{i+1}-u_{i}\parallel\leq\mu_{u}<\infty$.
Assumption 3
The deep neural network observable function $g(\cdot,\theta_{\tau})$ is Lipschitz continuous on the system state space with Lipschitz constant $\mu_{g}$.
Consider a DNN observable function $g(\cdot,\theta_{\tau}):\mathbb{R}^{n}\rightarrow\mathbb{R}^{r}$, let $\bar{\psi}^{h}=[\bar{\psi}_{1}^{h},\bar{\psi}_{2}^{h},\cdots,\bar{\psi}_{n_{h}}^{h}]^{T}\in\mathbb{R}^{n_{h}}$ be its last hidden layer and $\bar{\psi}^{o}=[\bar{\psi}_{1}^{o},\bar{\psi}_{2}^{o},\cdots,\bar{\psi}_{r}^{o}]^{T}\in\mathbb{R}^{r}$ denotes its output layer, where $\bar{\psi}_{i}^{h}:\mathbb{R}\rightarrow\mathbb{R}$ and $\bar{\psi}_{i}^{o}:\mathbb{R}\rightarrow\mathbb{R}$ are generally nonlinear function chosen by user. Note that here $n_{h}$ and $r$ denote the number of nodes of $\bar{\psi}^{h}$ and $\bar{\psi}^{o}$, respectively.
Assumption 4
(1): Given a Hilbert space $\mathcal{F}$, the Koopman operator $\mathcal{K}$ is bounded and continuous on $\mathcal{F}$; (2): $\bar{\psi}^{h}$ are selected from the orthonormal basis of $\mathcal{F}$, i.e., $[\bar{\psi}_{1}^{h},\bar{\psi}_{2}^{h},\cdots,\bar{\psi}_{\infty}^{h}]^{T}$ is an orthonormal basis of $\mathcal{F}$.
Remark 2
Since the DNN observable function $g(x,\theta_{\tau})$ is with known structure, one can make Assumption 4 hold by choosing proper activation functions for DNN’s different layers. We refer to [18] for more details about candidate functions like radial basis functions.
Lemma 2
Let $\mu\in\mathbb{R}$ be the empirical measurements associated with the observed states $x_{k}\in\mathbb{R}^{n}$. If Assumption 4 holds, then the operator $P_{n_{h}}^{\mu}$ converges strongly to the identity operator $I$ as $n_{h}$ goes to infinity.
Proof of Lemma 2 is referred to in [15].
Lemma 3
Let $P\coloneqq\begin{bmatrix}P_{n_{h}}^{\mu_{1}}&\mathbf{0}_{n_{h}\times m}\\ \mathbf{0}_{m\times n_{h}}&P_{m}^{\mu_{2}}\end{bmatrix}$ with $\mu_{1}\in\mathbb{R},\mu_{2}\in\mathbb{R}$ the empirical measurements associated to the observed system state $x_{k}\in\mathbb{R}^{n}$ and control input $u_{k}\in\mathbb{R}^{m}$, respectively. If Assumption 4 holds, then the approximated sequence of operators $\mathcal{K}_{D}P=P\mathcal{K}P$ converges to Koopman operator $\mathcal{K}$ strongly as $n_{h}$ goes to infinity.
Proof of Lemma 3 is given in [15].
Remark 3
Lemma 3 shows the convergence condition of the existing DKO method regarding the DNN observable function $g(\cdot,\theta_{\tau})$ ($n_{h}\rightarrow\infty$). It replaces the convergence condition of the EDMD method [17] ($r\rightarrow\infty$). This conclusion is used in the present work to reduce $\beta_{\tau}\geq r+m$ so that one can track the unknown NTVS based on $\mathcal{B}_{\tau}$ with small batch size.
Theorem 1
Let $g(x,\theta_{\tau})$ be a DNN with its last hidden layer containing $n_{h}$ nodes. If Assumptions 1-4 hold, then the estimation error $e_{k}$ in (22) is bounded by
$\lim_{n_{h}\rightarrow\infty}\sup\parallel e_{k}\parallel=(\parallel C_{\tau}A_{\tau}\parallel\mu_{g}+1)\mu_{x}+\parallel C_{\tau}B_{\tau}\parallel\mu_{u}+\max_{\bar{x}\in\mathcal{B}_{\tau}^{x}}\parallel\bar{x}-C_{\tau}g(\bar{x},\theta_{\tau})\parallel.$
The proof of Theorem 1 is given in [15].
Remark 4
Theorem 1 says that with certain assumptions hold, as $n_{h}\rightarrow\infty$, the estimation error of the proposed method depends on the minimization performance of (13) and $\mu_{x},\mu_{u}$ defined in Assumption 2. One way to further reduce the estimation error is to decrease the sampling interval.
Since one cannot construct a DNN with infinite parameters in practice, we introduce the following corollary based on Theorem 1, for which we need to make the following assumption.
Assumption 5
For any $A_{\tau}\in\mathbb{R}^{r\times r}$, $\parallel A_{\tau}\parallel<1$.
Remark 5
According to Lemma 1, the DNN observable function determines $\parallel A_{\tau}\parallel$. To make Assumption 5 hold, one can add an extra loss function to impose the maximum value of $\parallel A_{\tau}\parallel$ in (9).
Corollary 1
Recall $e_{k}$ in (22). If Assumptions 1-3 and 5 hold, then the upper bound of $\parallel e_{k}\parallel$ is determined by the minimization performance of (19) and $\mu_{x},\mu_{u}$ as $k$ goes to infinity.
The proof of Corollary 1 is given in [15].
4 Numerical Simulations
In this section, we test the proposed algorithm in two examples, one is a simple nonlinear time-varying system (NTVS), and the other one is a quadcopter. In both examples, the system state input data is observed with a fixed time interval of $0.1$s, and $\beta_{\tau}\equiv\beta$, where $\beta\geq r+m$ is an arbitrary positive integer, and $\tau=\lceil\frac{k-\beta_{0}}{\beta}\rceil$, where $k\in\mathbb{K}_{\tau}$.
4.1 A Simple NTVS
Consider the following dynamical system:
$\dot{x}_{t}=M_{t}\cos(x_{t}),$ (24)
where $x_{t}\in\mathbb{R}^{2}$ and $M_{t}$ is a time-varying matrix given by
$M_{t}=\begin{bmatrix}0&(1+\gamma t)\\ -(1+\gamma t)&0\end{bmatrix},$ (25)
where $\gamma$ is a constant determining how fast the dynamics change. For more details of the dynamics, see [14].
Experiment Setup. The DNN observable function ($g(\cdot,\theta):\mathbb{R}^{2}\rightarrow\mathbb{R}^{6}$) is built with a hidden layer ReLu() that contains $32$ nodes and an output layer Relu() consisting of $6$ nodes. The DNN training process is implemented by choosing optimizer Adam [19] with learning rate = 1e-3 and weight decay rate = 1e-4. Then $\mathcal{B}_{\tau}$ is observed with $\beta=10$ starting from $x_{0}=[1,0]^{T}$.
Remark 6
The activation function $ReLu()$ is selected for this example because by appropriately selecting $x_{0}$ and $\gamma$, one can ensure that $x_{t}\geq 0$. This choice results in $ReLu(x)=x$ which belongs to the category of linear radial basis functions [20].
Results Analysis. We test the proposed method in the NTVS in (24) for both slow and fast-changing scenarios setting $\gamma=0.8$ and $\gamma=6$, respectively. The trajectories estimated by time-varying DMD(TVDMD) and the proposed method are shown in Figs. 2, and the estimation errors of both methods are exhibited in Fig. 3. Here, we let $x_{k}\in\mathbb{R}^{2}$ denote the true system states observed from the unknown NTVS. Let $\tilde{x}_{k}\in\mathbb{R}^{2}$ denote the estimated states generated by the TVDMD method from [14]. Let $\hat{x}_{k}\in\mathbb{R}^{2}$ be the estimated states from the proposed method (DKTV), as given in (23). Consequently, $\tilde{e}_{k}=\parallel\tilde{x}_{k}-x_{k}\parallel$ and $e_{k}=\parallel\hat{x}_{k}-x_{k}\parallel$ denote the estimation errors induced by the TVDMD and DKTV methods, respectively.
(a) $\gamma=0.8$
(b) $\gamma=6$
Figure 2: System states trajectories: TVDMD (black) vs. DKTV (blue).
Figure 3: System states estimation errors: TVDMD (black) vs. DKTV (blue).
It can be observed that when the NTVS varies slowly, both methods can capture the dynamics of the NTVS with reasonable accuracy. In situations where the NTVS varies rapidly, the estimation errors of the TVDMD method increase correspondingly. On the contrary, the proposed method can achieve consistent performance, benefiting from its DNN observable function.
4.2 System States Prediction for Single Quadcopter
In this example, we implement the proposed algorithm to predict the system states of a quadcopter with a time-varying disturbance applied to its body. The dynamics of the quadcopter is described as
$\begin{pmatrix}\dot{p_{n}}\\ \dot{p_{e}}\\ \dot{z}\end{pmatrix}$ $=\begin{pmatrix}c_{\theta}c_{\psi}&s_{\phi}s_{\theta}c_{\psi}-c_{\phi}s_{\psi}&c_{\phi}s_{\theta}c_{\psi}+s_{\phi}s_{\psi}\\ c_{\theta}c_{\psi}&s_{\phi}s_{\theta}s_{\psi}+c_{\phi}c_{\psi}&c_{\phi}s_{\theta}s_{\psi}-s_{\phi}c_{\psi}\\ s_{\theta}&-s_{\phi}c_{\theta}&-c_{\phi}c\theta\end{pmatrix}\begin{pmatrix}\dot{x}\\ \dot{y}\\ \dot{z}\end{pmatrix},$
$\begin{pmatrix}\dot{\phi}\\ \dot{\theta}\\ \dot{\psi}\end{pmatrix}$ $=\begin{pmatrix}1&s_{\phi}t_{\theta}&c_{\phi}t_{\theta}\\ 0&c_{\phi}&-s_{\phi}\\ 0&\frac{s_{\phi}}{c_{\theta}}&\frac{c_{\phi}}{s_{\theta}}\end{pmatrix}\begin{pmatrix}p\\ q\\ r\end{pmatrix},$
$\begin{pmatrix}\ddot{x}\\ \ddot{y}\\ \ddot{z}\end{pmatrix}$ $=\begin{pmatrix}r\dot{y}-q\dot{z}\\ p\dot{z}-r\dot{x}\\ q\dot{x}-p\dot{y}\\ \end{pmatrix}+\frac{1}{m}(\begin{pmatrix}f_{x}\\ f_{y}\\ f_{z}\end{pmatrix}+w_{t}),$
$\begin{pmatrix}\dot{p}\\ \dot{q}\\ \dot{r}\end{pmatrix}$ $=\begin{pmatrix}\frac{J_{y}-J_{z}}{J_{x}}qr\\ \frac{J_{z}-J_{x}}{J_{y}}pr\\ \frac{J_{x}-J_{y}}{J_{z}}pq\\ \end{pmatrix}+\begin{pmatrix}\frac{1}{J_{x}}\tau_{\phi}\\ \frac{1}{J_{y}}\tau_{\theta}\\ \frac{1}{J_{z}}\tau_{\psi}\\ \end{pmatrix},$
where $s_{\theta}$, $s_{\phi}$, $s_{\psi}$, $c_{\theta}$, $c_{\phi}$, $c_{\psi}$, and $t_{\theta}$ denote $\sin(\theta)$, $\sin(\phi)$, $\sin(\psi)$, $\cos(\theta)$, $\cos(\phi)$, $\cos(\psi)$, and $\tan(\theta)$, respectively. $p_{n}$, $p_{e}$, and $z$ denote the inertial north position, the inertial east position, and the altitude, respectively. $\dot{x}$, $\dot{y}$, and $\dot{z}$ denote the velocity along the x-axis, y-axis, and z-axis, respectively. $\phi$, $\theta$, and $\psi$ denote the roll angle, pitch angle, and yaw angle, respectively, and $p$, $q$, and $r$ denote the roll rate, pitch rate, and yaw rate, respectively. $f_{x}$, $f_{y}$, and $f_{z}$ denote the total force applied to the quadcopter along the x-axis, y-axis, and z-axis, respectively, and $w_{t}\in\mathbb{R}^{3}$ denotes the time-varying disturbance. $\tau_{\phi}$, $\tau_{\theta}$ and $\tau_{\psi}$ denote the torque of roll, pitch, and yaw, respectively, while $J_{x}$, $J_{y}$, and $J_{z}$ denote the inertia along the x-axis, y-axis, and z-axis, respectively. For further details on the dynamics of the quadcopter, see [21].
Experiment Setup. In order to compare the performance of the proposed algorithm with that of the single neural network method, we let $x_{k}\in\mathbb{R}^{12}$ denote the true state of the system, $\bar{x}_{k}\in\mathbb{R}^{12}$ and $\hat{x}_{k}\in\mathbb{R}^{12}$ denote the estimated states of the single DNN and the proposed method, respectively, and $\bar{e}_{k}=\parallel\bar{x}_{k}-x_{k}\parallel$ and $e_{k}=\parallel\hat{x}_{k}-x_{k}\parallel$ denote the estimation errors of the single DNN and DKTV methods, respectively. The data batch $\mathcal{B}_{\tau}$ is collected with $\beta=30$ manually controlling the quadcopter from position $(0,0,0)$ (meter) to $(1,2,3)$ (meter) with the time-varying disturbance force $w_{t}\in\mathbb{R}^{3}$ applied on the quadcopter. $w_{t}$ is generated from the standard normal distribution. For the DKTV method, the DNN observable function $g(\cdot,\theta):\mathbb{R}^{12}\rightarrow\mathbb{R}^{16}$ is built with one hidden layer Gaussian function containing $64$ nodes and one output layer Relu() consisting of 16 nodes. For the single DNN method, DNN $N(\cdot,\nu):\mathbb{R}^{16}\rightarrow\mathbb{R}^{12}$ is built with the same structure as $g(\cdot,\theta)$. Then the DNN training process of both methods is implemented by choosing optimizer Adam [19] with learning rate = 1e-3 and weight decay rate = 1e-4. The optimal vector $\nu\in\mathbb{R}^{p}$ is found by
$\nu^{*}=\arg\min_{\nu\in\mathbb{R}^{p}}L(\nu)=\arg\min_{\nu\in\mathbb{R}^{p}}\parallel N(x_{k},u_{k},\nu)-x_{k+1}\parallel,$ (26)
which leads to $\bar{x}_{k+1}=N(x_{k},u_{k},\nu^{*})$.
Results Analysis. We first show the DNN training process of both methods with $\tau=1,\cdots,7$ in Fig. 4, where one epoch denotes one forward and one backward pass through the DNN and $L(\nu),L(\theta)$ are defined in (26) and (19), respectively. As shown, DKTV needs fewer training epochs to converge and maintains a much lower loss value during the training process than the single DNN method, as expected due to its double minimization of (12)-(13).
Figure 4: Training process: DNN (black) vs. DKTV (blue).
Then in Fig. 5, we demonstrate the algorithm performance by showing the system states estimation errors of both methods. As is shown, the proposed algorithm can provide better prediction performance compared with the Single DNN method.
Figure 5: System states estimation error: DNN (black) vs. DKTV(blue).
Finally, in Fig. 6, Theorem 1 is validated by showing that the estimation errors of the proposed algorithm are reduced by increasing the number of nodes of its hidden layer.
Figure 6: The estimation errors of DKTV with different DNN observable functions.
4.3 Optimal Control based on the Proposed Deep Koopman Learning
In this subsection, we design the optimal controller based on the proposed method for a classic cartpole example from [22], where we make the coefficient of friction of cart on track ($\mu_{c}(t)$) time-varying during the simulation. And we choose the well-studied model predictive control (MPC) method to demonstrate the optimal control application based on the proposed algorithm, of which the optimization problem is formulated as
{mini}
[…]
$J(u_{i},x_{i})=$ $\sum_{i=k_{\tau}}^{k_{\tau}+l-1}(\bar{g}(x_{i},\theta_{\tau})^{T}\tilde{Q}_{\tau}\bar{g}(x_{i},\theta_{\tau})+u_{i}^{T}Ru_{i})+g(x_{l},\theta_{\tau})^{T}{\tilde{Q}_{l}}g(x_{l},\theta_{\tau}),$
and $l\leq\beta_{\tau}$ is the time horizon; $i$ is the time-index along the defined time horizon; $x_{i}\in\mathbb{R}^{n}$ and $u_{i}\in\mathbb{R}^{m}$ are the system state and control input at time step $i$, respectively; $\bar{g}(x_{i},\theta_{\tau})=g(x_{i},\theta_{\tau})-g(x^{*},\theta_{\tau})$ with $x^{*}$ denoting the goal state; $x_{l}$ is the terminal state; $\mathcal{X}$ denotes the state constraint and $\mathcal{U}$ denotes the control input constraint; $\tilde{Q}_{\tau}=C_{\tau}^{T}QC_{\tau}\in\mathbb{R}^{r\times r}$, $Q\in\mathbb{R}^{n\times n},R\in\mathbb{R}^{m\times m}$ are positive definite matrices.
4.3.1 Time-varying Cartpole Balance
In this example, the physic governing equation of the cartpole is described as
$\ddot{x}_{t}$ $=\frac{F_{t}+ml(\dot{\bar{\theta}}_{t}^{2}\sin{\bar{\theta}_{t}}-\ddot{\bar{\theta}}_{t}\cos{\bar{\theta}_{t}})-\mu_{t}^{c}\text{sgn}{(\dot{x}_{t}})}{m_{c}+m}$ (27)
$\ddot{\bar{\theta}}_{t}$ $=\frac{\cos{\theta_{t}}[-F_{t}-ml\dot{\bar{\theta}}_{t}^{2}\sin{\theta_{t}}+\mu_{t}^{c}\text{sgn}{(\dot{x}_{t}})]/(m_{c}+m)}{l[\frac{4}{3}-(m\cos^{2}{\bar{\theta}_{t}})/(m_{c}+m)]}+\frac{g\sin{\theta_{t}}-\mu_{p}\dot{\bar{\theta}}/{ml}}{l[\frac{4}{3}-(m\cos^{2}{\bar{\theta}_{t}})/(m_{c}+m)]},$
where $\dot{\mu}_{t}^{c}=0.3\cos(t)$ with $\mu_{0}^{c}=0.0005$ is the time-varying coefficient of friction of cart on track; $x_{t}$ denotes the distance of the cartpole moves from the initial position; $\bar{\theta}_{t}$ denotes the angle from the up position; $\dot{x}_{t},\dot{\bar{\theta}}_{t}$ denote the x-axis velocity and the angular velocity respectively; $F_{t}$ denotes the continuous control input applied to the center of the mass of the cart at time step $t$; $g=-9.8m/s^{2}$ is the gravity acceleration; $m_{c}=1.0kg,m=0.1kg$ are the mass of the cart and the pole, respectively; $l=0.5m$ is the length of the pole; $\mu_{p}=0.000002$ is the coefficient of the friction of the pole on cart.
For the simulation, we first build the DNN ($g(\cdot,\theta):\mathbb{R}^{4}\rightarrow\mathbb{R}^{6}$) and set $\beta=12$; then the $\mathcal{B}_{\tau}$ is generated by observing the dynamics in (27) with fixed time interval $t_{k_{\tau}+1}-t_{k_{\tau}}\equiv 0.1s$. The approximated dynamics is directly applied to design the MPC controller to keep the cartpole balanced at the up position with $\bar{\theta}=0,\dot{\bar{\theta}}=0$.
Results Analysis: Fig. 7 shows the trajectory of the time-varying cartpole under the MPC control based on the proposed method. From the result, the cartpole is able to keep the desired up position under the DKTV-based MPC control even though $\mu_{k}^{c}$ increases from $0.0005$ to $14.288$.
Figure 7: MPC based on the Approximated Dynamics
Remark 7
In this subsection, we show the MPC design based on approximated linear dynamics from the proposed algorithm. Compared to the existing MPC method for the NTVS proposed in [23, 24, 25], the proposed method requires neither the priori knowledge of the system dynamics nor the distribution information of the system disturbance.
5 Conclusions
This paper presents a data-driven approach to approximate the dynamics of a nonlinear time-varying system (NTVS) by a linear time-varying system (LTVS), which is resulted from the Koopman operator and deep neural networks. The proposed algorithm is able to approximate well an NTVS by iterative updating its linear approximation as validated by a simple NTVS system in two dimensions. Moreover, controllers developed based on linear approximation perform well in controlling a quadrotor with complex dynamics. Future work includes employing the proposed method to achieve adaptive autonomy in [26, 27] and to serve as robot’s dynamics estimation in human-robot teaming [28, 29].